# Heterogeneous Low-Rank Approximation for Federated Fine-tuning of On-Device Foundation Models

### 12 Jan 2024 v1

@@ 3.4 Why not Simply Reconstruct First,
[…]
@@
Table 3: Final RougeL score for Reddit text summarization and perplexity for multi-session chat for different federated fine-tuning methods. The blue text indicates the ratio of trained number of parameters compared to the full-fine tuning case. HetLoRA outperforms both HomLoRA and Recon+SVD method, but slightly underperforms the full-fine tuning case. However, compared to full fine-tuning the number of trained parameter is significantly smaller.
@@
PaLM-XXS PALM-XS PaLM-XXS PALM-XS
@@ 4 Experiments
In this section, we present results for HetLoRA and its baselines in terms of the performance on training speed, computation/communication efficiency, and final achieved performance. First, we show the performance of homogeneous LoRA to show how LoRA in general performs for low and high rank values. Second, we demonstrate HetLoRA’s performance for different $r_{\text{min}}$ and $r_{\text{max}}$ values comparing them with full-fine tuning, homogeneous LoRA, and the reconstruction-first method elaborated in Section 3.4. We also conduct an ablation study on HetLoRA with varying decay factor $\gamma$ for the rank self-pruning step.The rank distribution across clients for HetLoRA, unless mentioned otherwise, is set to a truncated power-law distribution with $\alpha=0.1$ in the range between $[r_{\text{min}},r_{\text{max}}]$ (inclusively), where the small $\alpha$ value makes the distribution skewed towards smaller ranks. All experiments were ran with 3 different random seeds and their average is shown along with the standard deviation.
Model. We use the transformer-based model PaLM2 Google (2023) of size XXS and XS for our experiments which we consider applicable to fit in to the category of ODFMs given their smaller sizes (340M-1B number of parameters) compared to standard FMs. The LoRA modules are applied to only the self-attention layers as proposed in the original LoRA paper Hu et al. (2021), and their relative number of parameters compared to the original model are shown in Table 2.
@@ 4.1 Experiment Results
PaLM-XXS PALM-XS PaLM-XXS PALM-XS
$\gamma=1$ $92.17(\pm 0.08)$ $91.95(\pm 0.03)$ $55.07(\pm 0.81)$ $40.92(\pm 0.58)$
$\gamma=0.99$ $\mathbf{94.23}(\pm 0.03)$ $\mathbf{94.41}(\pm 0.05)$ $\mathbf{53.93}(\pm 1.57)$ $\mathbf{38.76}(\pm 0.52)$
$\gamma=0.95$ $89.62(\pm 1.33)$ $83.19(\pm 1.70)$ $71.10(\pm 1.39)$ $46.39(\pm 0.87)$
$\gamma=0.85$ $60.31(\pm 3.04)$ $53.28(\pm 2.47)$ $120.72(\pm 10.93)$ $59.67(\pm 1.98)$
Heterogeneous LoRA compared to Baselines. Finally, we compare our proposed HetLoRA with other baselines in Table 3 and Fig. 5. We see that HetLoRA with $r_{\text{min}}=5$ and $r_{\text{max}}=50$ achieves faster training as well as better performance than homogeneous LoRA cases with both edge cases of the ranks $r\in\{5,50\}$ and reconostruction+SVD which was explained in Section 3.4. This implies that HetLoRA is not only practical in the sense that clients are allowed to have their own rank values, it can also outperform the limited case of homogeneous LoRA where all clients have $r=r_{\text{min}}$ or the impractical case where all clients have $r=r_{\text{max}}$. We also observe that HetLoRA achieves slightly lower performance than full fine-tuning. However, as shown in the blue text in Table 3 that shows the number of trained parameters compared to the full-fine tuning case, full fine-tuning requires to train a much larger number of parameters compared to HetLoRA, making it infeasible to train with ODFMs in practice. We also show in Fig. 6 that to achieve the targeted performance for both Reddit and MSC task, HetLoRA requires significantly less number of parameters to be trained and communicated compared to full-fine tuning. Although for Reddit, HomLoRA has a slightly less number of parameters to be trained, the final achieved RougeL is outperformed by HetLoRA as shown in Table 3.
@@
Figure 5: Comparison of the performance across homogeneous LoRA, heterogeneous LoRA, and full-fine tuning. Heterogeneous LoRA achieves better performance than homogeneous LoRA with fewer number of communication rounds.

### 20 Feb 2024 v2

@@ 3.4 Why not Simply Reconstruct First, then Redistribute the LoRA modules?
Let us consider a simple case where there are 2 clients with heterogeneous rank lora modules $\mathbf{B}_{1}\in\mathbb{R}^{d\times 1},\mathbf{A}_{1}\in\mathbb{R}^{1\times l}$ and $\mathbf{B}_{2}\in\mathbb{R}^{d\times 1},\mathbf{A}_{2}\in\mathbb{R}^{2\times l}$ respectively for client 1 and client 2 where the former has rank $1$ and latter has rank $2$. We set the notation for the LoRA modules’ $i^{th}$ row and $j^{th}$ column value for $\mathbf{B}_{k}$ and $\mathbf{A}_{k}$ as $b_{k,ij}$ and $a_{k,ij}$ respectively. Then with $d=3,\leavevmode\nobreak\ l=2$, when we reconstruct each of the LoRA modules first and then aggregate the full model we have its $i^{th}$ row and $j^{th}$ column as $(\sum_{k=1}^{2}b_{k,i0}a_{k,0j})+b_{2,i1}a_{2,1j}$ and aggregating the LoRA modules first and then reconstructing the model has the full model’s $i^{th}$ row and $j^{th}$ column as $(\sum_{k=1}^{2}b_{k,i0})(\sum_{k=1}^{2}a_{k,0j})+b_{2,i1}a_{2,1j}$.
@@
Table 3: Final RougeL score for Reddit text summarization and perplexity for multi-session chat for different federated fine-tuning methods. The blue text indicates the ratio of trained number of parameters compared to the full fine-tuning case. HetLoRA outperforms both HomLoRA and Recon+SVD method, but slightly underperforms the full fine-tuning case. However, compared to full fine-tuning the number of trained parameter is significantly smaller.
@@
PaLM 2-XXS PaLM 2-XS PaLM 2-XXS PaLM 2-XS
@@ 4 Experiments
In this section, we present results for HetLoRA and its baselines in terms of the performance on training speed, computation/communication efficiency, and final achieved performance. First, we show the performance of homogeneous LoRA to show how LoRA in general performs for low and high rank values. Second, we demonstrate HetLoRA’s performance for different $r_{\text{min}}$ and $r_{\text{max}}$ values comparing them with full fine-tuning, homogeneous LoRA, and the reconstruction-first method elaborated in Section 3.4. We also conduct an ablation study on HetLoRA with varying decay factor $\gamma$ for the rank self-pruning step.The rank distribution across clients for HetLoRA, unless mentioned otherwise, is set to a truncated power-law distribution with $\alpha=0.1$ in the range between $[r_{\text{min}},r_{\text{max}}]$ (inclusively), where the small $\alpha$ value makes the distribution skewed towards smaller ranks. All experiments were ran with 3 different random seeds and their average is shown along with the standard deviation.
Model. We use the Transformer-based language model PaLM 2 Google (2023) of size XXS and XS for our experiments which are lightweight enough to fit in to the category of ODFMs Google DeepMind (2023) compared to standard FMs. The LoRA modules are applied to only the self-attention layers as proposed in the original LoRA paper Hu et al. (2021), and their relative number of parameters compared to the original model are shown in Table 2.
@@ 4.1 Experiment Results
PaLM 2-XXS PaLM 2-XS PaLM 2-XXS PaLM 2-XS
HetLoRA, $\gamma=1$ $92.17\leavevmode\nobreak\ (\pm 0.08)$ $91.95\leavevmode\nobreak\ (\pm 0.03)$ $55.07\leavevmode\nobreak\ (\pm 0.81)$ $40.92\leavevmode\nobreak\ (\pm 0.58)$
HetLoRA, $\gamma=0.99$ $\mathbf{94.23}\leavevmode\nobreak\ (\pm 0.03)$ $\mathbf{94.41}\leavevmode\nobreak\ (\pm 0.05)$ $\mathbf{53.93}\leavevmode\nobreak\ (\pm 1.57)$ $\mathbf{38.76}\leavevmode\nobreak\ (\pm 0.52)$
HetLoRA, $\gamma=0.95$ $89.62\leavevmode\nobreak\ (\pm 1.33)$ $83.19\leavevmode\nobreak\ (\pm 1.70)$ $71.10\leavevmode\nobreak\ (\pm 1.39)$ $46.39\leavevmode\nobreak\ (\pm 0.87)$
HetLoRA, $\gamma=0.85$ $60.31\leavevmode\nobreak\ (\pm 3.04)$ $53.28\leavevmode\nobreak\ (\pm 2.47)$ $120.72\leavevmode\nobreak\ (\pm 10.93)$ $59.67\leavevmode\nobreak\ (\pm 1.98)$
Heterogeneous LoRA compared to Baselines. Finally, we compare our proposed HetLoRA with other baselines in Table 3 and Fig. 5. We see that HetLoRA with $r_{\text{min}}=5$ and $r_{\text{max}}=50$ achieves faster training as well as better performance than homogeneous LoRA cases with both edge cases of the ranks $r\in\{5,50\}$ and reconostruction+SVD which was explained in Section 3.4. This implies that HetLoRA is not only practical in the sense that clients are allowed to have their own rank values, it can also outperform the limited case of homogeneous LoRA where all clients have $r=r_{\text{min}}$ or the impractical case where all clients have $r=r_{\text{max}}$. We also observe that HetLoRA achieves slightly lower performance than full fine-tuning. However, as shown in the blue text in Table 3 that shows the number of trained parameters compared to the full fine-tuning case, full fine-tuning requires to train a much larger number of parameters compared to HetLoRA, making it infeasible to train with ODFMs in practice. We also show in Fig. 6 that to achieve the targeted performance for both Reddit and MSC task, HetLoRA requires significantly less number of parameters to be trained and communicated compared to full fine-tuning. Although for Reddit, HomLoRA has a slightly less number of parameters to be trained, the final achieved RougeL is outperformed by HetLoRA as shown in Table 3.
@@
Figure 5: Comparison of the performance across homogeneous LoRA, heterogeneous LoRA, and full fine-tuning. Heterogeneous LoRA achieves better performance than homogeneous LoRA with fewer number of communication rounds.
@@ 5 Discussions and Concluding Remarks
6 Limitations
In this work, we address tackling system and data heterogeneity in federated fine-tuning of on-device foundation models. Our work is motivated by clients being able to carry different ranks for the LoRA fine-tuning method depending on their available resources, and thus exploiting this characteristic to improve federated fine-tuning with heterogeneous LoRA. However, our work assumes that the rank distribution across clients (which is analogous to how system resources are distributed across clients) is independent to the data distribution. There can be scenarios in which this is not necessarily the case where the rank and data distribution can be correlated. For instance, more affluent populations can have better off devices with larger resource capacity, and may have data distributions different to that of less affluent populations. Such correlation should be explored for future work to better understand the implications of heterogenoeus LoRA.